x divided by 7 = 9 Solve for x
Answer by oberobic(2304)

x/7 = 9
.
multiply both sides of the equation by 7
.
7*(x/7) = 7*9
.
on the left-hand side, the 7's cancel, leaving you with x
on the right-hand side, 7*9 = 63
.
x = 63